# Dressed to Gamble: How Poker Drives the Dynamics of Wearables and Visits on Decentraland’s Social Virtual World

### v2

@@ Abstract
Decentraland is a blockchain-based social virtual world touted to be a creative space owned by its community. In it, users can publish wearables used to customize avatars, which can be then sold or given away via blockchain transfers. Decentral Games (DG), a single project owning prominent in-world casinos, has by far created the most wearables, necessary to earn cryptocurrency in its flagship game ICE Poker. Herein, we present a comprehensive study on how DG and ICE Poker influence the dynamics of wearables and in-world visits in Decentraland. To this end, we analyzed 5.9 million wearable transfers made on the Polygon blockchain (and related sales) over a two-year period, and 677 million log events of in-world user positions in an overlapping 10-month period. We found that the platform-wise number of transfers and sales monetary value of wearables were disproportionally related to DG, and that its two ICE Poker casinos (less than 0.1% of the world map) represented a very large average share of daily unique visitors (33%) and time spent in the virtual world (20%). Despite several alternative in-world economic and artistic initiatives in Decentraland, some of which have attracted much attention from the general public, a single third-party online poker game appears to be the main driver of the analyzed dynamics. Our work thus contributes to the current understanding of user behavior in social virtual worlds, and it is among the first to study the emerging phenomenon of blockchain-based online gambling in virtual reality spaces.
keywords:
@@
social virtual worlds, Metaverse, online gambling, NFT, blockchain
1 Introduction
@@
In recent years, there has been much discussion around the Metaverse, an envisioned fully immersive iteration of the Internet, composed of a shared collection of interoperable 3D virtual worlds. The Metaverse remains a vision that is in its early stages Cheng et al. (2022), but many social virtual worlds usually associated with it thrived during the Covid-19 pandemic Kerdvibulvech (2022). In particular, interest surged in the wake of Facebook’s rebranding into Meta in late 2021, which signaled the company’s quest to realize such an ambitious vision Kraus et al. (2022). In these virtual worlds, interaction with the environment and with other people is done primarily through an avatar, an embodied representation of the user. Further, avatar customization (e.g., via virtual wearables) is a significant in-world factor of purchase intent, as well as one of the main pleasures for users Bleize & Antheunis (2019). However, it has been argued Zhou et al. (2018) that many of the ownership policies for in-world digital goods —such as those used in avatar customization— are disadvantageous to users. This is primarily due to the imbalance in rights between the users and the proprietors of the virtual world platforms. For example, traditionally digital assets are not transferable independently of the platform; hence, these assets remain locked-in inside the virtual world or are forever lost when the underlying platform ceases to function.
@@
In response, blockchain technology has been heralded by many tech enthusiasts as a decentralized and transparent approach to digital property in virtual worlds Cannavo & Lamberti (2020). Arguably, Decentraland is the earliest and most popular blockchain-based virtual world, touted to be the first one “built, governed, and owned by its users” in the project’s official website. In Decentraland, users can transfer in-world land, usernames, and avatar wearables, either as part of a sale or for free. All of these transfers are recorded on the blockchain Ethereum, or in the case of wearables on Polygon, a compatible and more scalable blockchain. Hence, there is a trove of public data regarding the dynamics of this social virtual world.
@@
Based on these data, Trujillo and Bacciu 2023 were among the first to conduct a quantitative study of Decentraland’s wearables. In their analysis, it emerged that a single account —related to the project Decentral Games (DG)— was by far the one with most published wearables. DG manages a few in-world virtual casinos, and possession of DG wearables is a requirement to earn money (in the form of cryptocurrency) by playing ICE Poker, the main project’s attraction (depicted in Figure 1). In addition, owners can delegate their DG wearables to other users, with whom earnings are split based on the wearable’s level, which can be also upgraded for a fee. Hence, DG wearables transcend the mere aesthetic in-world function of virtual garments, with ownership being enticing to users of Decentraland due to their earning potential.
Figure 1: Screenshot of an ICE Poker game session inside a Decentral Games casino on Decentraland. Note the avatar wearing the sleeveless shirt with the name of the game, which allows the player to earn cryptocurrency, as well as the croupier being based on the Internet meme Pepe the Frog.
@@
Therefore, intrigued by both the peculiar behavior of the aforementioned account managed by DG and the economic incentive to play in its casinos, herein we delve into how DG and ICE Poker influence the dynamics of ownership transfers of wearables and in-world visits in Decentraland. In particular, we seek to answer the following research questions (RQs):
@@
To answer RQ1, we measure the share of wearables designed by DG in the overall dynamics of transferring ownership of Decetranland’s wearables in its different forms, including creation, sale, and destruction, as well as how the upgrading mechanism implemented by DG affects such transfers. To answer RQ2, we quantify the visits —both in time and number of visitors— made on parcels in which ICE Poker casinos are located, compared to the rest of the in-world land. We also consider and describe the time series characteristics of these visits on a daily basis.
1.1 Contributions and significance
@@
Our main contribution is the first quantitative analysis regarding the influence that online gambling has had on Decentraland’s social virtual world. Thanks to its trailblazing development based on blockchain and tokenomics, Decentraland gives a glimpse of how new kinds of immersive social media might work on the so-called Metaverse. Therefore, studying the use (and potential abuse) of the platform’s affordances and economic incentives in the form of online gambling is important to better shape the future mechanisms and legal framework of social virtual worlds.
2 Background
@@
In this section we briefly describe the main mechanisms, jargon, and events of both Decentraland and Decentral Games referred to in the related work, analyses, and discussion thereof.
2.1 Blockchain technologies
@@
Blockchains are decentralized, immutable ledgers that facilitate secure and transparent recording of transactions across a network of computers. They are the technological backbone of cryptocurrencies such as Bitcoin (BTCNakamoto (2008) and Ether (ETHWoodet al. (2014), digital currencies that enable financial transactions without the need for a central authority like a bank. Blockchains also provide the underlying mechanisms for non-fungible tokens (NFTs), unique identifiers often used to represent ownership of digital objects (e.g, art and collectibles) Nadini et al. (2021). Unlike fungible cryptocurrencies, NFTs are unique and non-interchangeable. Furthermore, blockchains enable decentralized autonomous organizations (DAOs), a form of token-based governance without a central authority, allowing for distributed decision-making and management Hassan & De Filippi (2021).
The Ethereum blockchain Woodet al. (2014), in particular, is a versatile platform that supports all of the aforementioned applications, utilizing smart contracts —self-executing programs that automatically enforce the terms of an agreement. Decentraland’s features are based on the Ethereum blockchain, except for transactions concerning wearables; these are instead based on Polygon Kanani et al. (2021), a higher layer scaling solution with its own cryptocurrency (MATIC); it is compatible with Ethereum and is designed to improve transaction speeds and reduce costs while maintaining security and decentralization.
2.2 Decentraland
Publicly released in February 2020, Decentraland’s social virtual world revolves around the so-called tokenomics, i.e., a decentralized economy based on the exchange of assets via cryptographic tokens on the blockchain Lo & Medda (2020). On Ethereum and Polygon, these exchanges are made from one address to another address via the aforementioned smart contracts, which set the automatic rules of the transactions. These addresses identify either an account or a smart contract that can participate on the blockchain.
@@ 2.2 Decentraland
Decentraland has its own cryptocurrency, called MANA, with which users can buy in-world digital assets, such as parcels of land and wearables. The proof of ownership of these assets is in the form of NFTs. Possession of MANA is also used as the basis of voting power for platform decisions via the decentralized autonomous organization of Decentraland Goldberg & Schär (2023).
@@
The world (see Figure 2) is composed of 90,601 square land parcels distributed on a 301$\times$301 grid, whose coordinates range in both axes from -150 to 150. Land is divided into two main types: public and privately owned. Public land is in turn divided into roads (10.4% of the world map) and plazas (4%), and cannot be sold. Before the public release of the project, 37.4% of the world map was allocated for districts, i.e., groups of adjacent parcels of privately owned land that share similar user interests. District parcels were assigned via an auctioning process to different groups, which respectively manage the land within their boundaries. Among the biggest districts we find, for example, the gambling-focused Vegas City and the shopping-oriented Fashion Street.
@@
Figure 2: Decentraland world map, where the biggest districts are indicated in violet and roads in gray.Vegas City district is home to Diamond Hands City and The Stronghold (both highlighted in yellow), ICE Poker casinos owned by Decentral Games. They cover only a minuscule surface of the map, accounting for less than 0.1%.
@@
Decentraland wearables are virtual garments, accessories or full body costumes (called skins) that can be used to clothe one’s avatar or completely change its appearance. Wearables are grouped into collections of one or more items. Each item represents a distinct design from which one or more NFTs can be minted, i.e, a new unique token ID representing the asset is created and assigned. Mints are registered on the blockchain as transfers from a special null address. In addition, a given item has a rarity limit (from a predefined list) that sets its maximum supply of tokens —e.g., a unique item can be minted only once, while a common item 100,000 times. To publish a wearable collection, creators pay a fee and submit their designs to the approval committee appointed by the DAO. Once approved, creators can sell their wearables via Decentraland’s Marketplace (for a fee that goes to the DAO), or they can mint and transfer the respective NFTs to a blockchain address (their own or not), either as a gift or to be sold on a different marketplace.
Originally, in the first version of wearables, these were created via a contest with approved designers and minted on Ethereum. However, this proved to be too limiting and prohibitive due to the then high transaction fees on that blockchain. Therefore, the wearables system was changed to the second and current version described above, now open to all and based on the less costly and more scalable Polygon blockchain. For the launch of both versions, a fashion event was held to showcase this creative endeavour and promote the newly published wearables; these events would be the precursors of more elaborate fashion initiatives. Arguably, the most important subsequent initiative was the first edition of the Metaverse Fashion Week in March 2022, a series of in-world social events that involved renowned fashion brands such as Tommy Hilfiger and Dolce & Gabbana.
2.3 Decentral Games
The project DG started in 2019, a year prior to Decentraland’s public release, and has played a part in many initiatives in the virtual world ever since. In fact, DG established one of the first in-world casinos, with games such as roulette or poker, and it was one of the few creators involved in the short-lived first version of Decentraland wearables. DG also established its own DAO, as well as its utility token and namesake DG, which grants access to rewards and voting power in the project’s governance.
@@ 2.3 Decentral Games
In October 2021, DG released ICE Poker, touted to be a Metaverse Play-to-Earn (P2E) game, i.e., one in which users are able to earn cryptocurrency simply by playing. In preparation, DG also launched the cryptocurrency ICE to function as the in-game currency for rewards and minting wearables. The following month, DG also published several exclusive ICE-themed wearables that granted attractive bonuses while playing the game, based on their ranking level. These were made available for minting in three batches to users that had a certain amount of DG and ICE.
@@
Inside Decentraland, there are two DG virtual casinos dedicated to ICE Poker, both located in the Vegas District. The first one and more popular is called The Stronghold; the second one, Diamond Hands City, is exclusive to accounts that possess ICE-themed wearables with the highest rank. Outside Decentraland, DG wearables also give access to a related but separate mobile and web game released on September 2022, now called ICE Poker Arcade. In this version of the game, instead of cryptocurrency players earn electronic gadgets, DG merchandise or virtual assets such as wearables.
@@
The level of access and/or reward provided by DG wearables varies by item, as illustrated by the examples in Figure 3; the more expensive the better. Once inside either game mode, players can wear up to five wearables to augment their rewards. For certain ICE-themed wearables, it is also possible to upgrade their ranking level (ranging from 1 to 5) to gain better rewards by paying a smaller fee compared to buying one. Additionally, all DG wearables can be delegated to other users, who can then play and split their earnings with the owner of the wearable, with the latter retaining the rights over the asset on the blockchain and being able of revoking the delegation at any moment.
@@
Figure 3: Examples of wearables created by Decentral Games: a) Venetian Mask (ICE Rank 1), which gives 1–7% bonus in ICE Poker; b) DG Suited Tracksuit Bottom, which can be used to play ICE Poker Arcade Mode; and c) Paludamentum (ICE Rank 5), which gives 35–45% bonus in ICE Poker.
3 Related Work
@@
While we are still in the early stages of Metaverse development, numerous independent virtual worlds have already sprung up over the past few decades. Notable examples include Second Life, Meta Horizon Worlds, and massively multiplayer online (MMO) games such as World of Warcraft, Minecraft, Roblox and Fortnite. However, these virtual environments have faced significant criticism regarding the governance and ownership of in-world assets. In Second Life, for instance, there has been a perceived emergence of a new form of virtual feudalism Grimmelmann (2009), while Roblox has been criticised for profiting from child labor Parkin (2022) and for allowing harmful content and games with gambling-like mechanisms in its user-generated virtual worlds Kou & Gui (2023). World of Warcraft has grappled with conflicts between the game’s publisher and players over agency and ownership of in-game assets Glas (2010). Additionally, there has been a general separation between content and platform ownership in virtual worlds, often to the detriment of users, whose rights over bought or created in-platform content are usually restricted Zhou et al. (2018).
In response to these challenges, a wave of blockchain-based virtual experiences has emerged Cannavo & Lamberti (2020); Huynh-The et al. (2023). Virtual worlds such as Decentraland Ordano et al. (2017), The Sandbox Madrid et al. (2020), Axie Infinity Nguyen et al. (2021), Upland Honigstein et al. (2019), and Illuvium Warwick & Warwick (2023) all attempt to mitigate the aforementioned issues by leveraging the main strengths of blockchain technology (transparency, security, and decentralization) for various purposes. These include verifying the digital ownership of virtual assets, enabling cryptocurrency-based economies, and implementing decentralized governance systems.
@@
In particular, Decentraland constitutes a standout example, being the earliest and most popular virtual world based on blockchain. It has reached a wide audience and gone through many changes over several years, hence it offers a trove of publicly available data for the study of emerging phenomena in these novel asset-oriented social media. However, existing research on Decentraland has primarily focused on landholding assets Dowling (2022); Goldberg et al. (2021); Guidi & Michienzi (2022) or examined the platform’s tokens within the broader context of the NFT ecosystem Nadini et al. (2021). There is a lack of studies regarding the other main kinds of digital assets available in Decentraland, namely usernames and wearables for avatars, with the latter having attracted the attention of the fashion industry Gonzalez (2022).
3.1 Fashion in Virtual Worlds

### v3

@@ Abstract
Decentraland is a blockchain-based social virtual world where users can publish and sell wearables to customize avatars. In it, the third-party Decentral Games (DG) allows players of its flagship game ICE Poker to earn cryptocurrency only if they possess certain wearables. Herein, we present a comprehensive study on how DG and its game influence the dynamics of wearables and in-world visits in Decentraland. To this end, we analyzed 5.9 million wearable transfers made on the Polygon blockchain (and related sales) over a two-year period, and 677 million log events of in-world user positions in an overlapping 10-month period. We found that these activities are disproportionately related to DG, with its ICE Poker casinos (less than 0.1% of the world map) representing a remarkable average share of daily unique visitors (33%) and time spent in the virtual world (20%). Despite several alternative initiatives within Decentraland, ICE Poker appears to drive user activity on the platform. Our work thus contributes to the understanding of how play-to-earn games influence user behavior in social virtual worlds, and it is among the first to study the emerging phenomenon of virtual blockchain-based gambling.
Amaury TRUJILLO
Istituto di Informatica e Telematica, Consiglio Nazionale delle Ricerche (IIT-CNR)
Pisa, Italy
amaury.trujillo@iit.cnr.it
[Uncaptioned image] $0$-0001-6227-0944
(Corresponding Author)
Clara BACCIU
Istituto di Informatica e Telematica, Consiglio Nazionale delle Ricerche (IIT-CNR)
Pisa, Italy
clara.bacciu@iit.cnr.it
[Uncaptioned image] $0$-0002-8711-7812
Matteo ABRATE
Istituto di Informatica e Telematica, Consiglio Nazionale delle Ricerche (IIT-CNR)
Pisa, Italy
matteo.abrate@iit.cnr.it
[Uncaptioned image] 0009-0004-6809-3570
@@
Keywords— social virtual worlds, play-to-earn, online gambling
I.   Introduction
@@
In recent years, there has been much discussion around the Metaverse, an envisioned fully immersive iteration of the Internet, composed of a shared collection of interoperable 3D virtual worlds. The Metaverse remains a vision in its early stages, but many social virtual worlds usually associated with it thrived during the Covid-19 pandemic [33]. In these worlds, interaction with the environment and with other people is done primarily through an avatar, an embodied representation of the user, with avatar customization being a significant in-world factor of purchase intent and one of the main pleasures for users [4]. However, it has been argued [58] that many of the ownership policies for in-world digital goods —such as those used in avatar customization— are disadvantageous to users. This is primarily due to the imbalance in rights between the users and platform proprietors. Traditionally, digital assets are not transferable independently of the platform; hence, these remain locked inside the virtual world or are forever lost when the underlying platform ceases to function.
@@
In response, blockchain technology has been heralded by many tech enthusiasts as a decentralized and transparent approach to digital property in virtual worlds [9]. Arguably, Decentraland is the earliest and most popular blockchain-based virtual world, touted to be the first one “built, governed, and owned by its users” in its official website. In Decentraland, users can transfer in-world land, usernames, and avatar wearables, either as part of a sale or for free. All these transfers are recorded on the Ethereum blockchain, or, in the case of wearables, on Polygon, a compatible and more scalable blockchain.
@@
Based on these data, Trujillo and Bacciu [51] were among the first to conduct a quantitative study of Decentraland’s wearables. They noticed that a single account —related to the project Decentral Games (DG)— had by far the most published wearables. DG manages a few in-world casinos, and possession of DG wearables is required to earn money (in the form of cryptocurrency) by playing ICE Poker, its main attraction (depicted in Figure I). Owners can also delegate their DG wearables to others, with whom earnings are split based on the wearable’s level. Hence, DG wearables transcend the mere aesthetic in-world function of virtual garments, with ownership being enticing due to their earning potential.
@@
Herein we thus delve into how DG and ICE Poker influence the dynamics of wearables’ ownership and in-world visits in Decentraland. In particular, we seek to answer the following research questions (RQs):
@@
To answer RQ1, we measure the share of wearables designed by DG in the overall dynamics of ownership transfers in different forms, including creation, sale, and destruction, as well as how the upgrading mechanism implemented by DG affects such transfers. To answer RQ2, we analyze the visits —both in time spent and number of visitors— made on parcels in which ICE Poker casinos are located, compared to the rest of the in-world land.
Figure I: Screenshot of an ICE Poker game session inside a Decentral Games casino.
@@
Our main contribution is therefore a quantitative analysis regarding the influence that online gambling has had on Decentraland’s social virtual world. In fact, integrating earning opportunities into virtual spaces has already resulted in serious repercussions in the real world[26, 41], emphasizing how virtual worlds can be both digital (in their essence) and real (in their effects) [7]. Thus, studying the use (and potential abuse) of Decentraland’s affordances and economic incentives in the form of online gambling is important to better shape the future mechanisms and legal framework of social virtual worlds and the Metaverse in general.
II.   Background
@@
In this section we briefly describe the main mechanisms, jargon, and events of both Decentraland and Decentral Games referred to in subsequent sections.
A.   Blockchain technologies
@@
Blockchains are decentralized, immutable ledgers that facilitate secure and transparent recording of transactions across a network of computers. Decentraland is based on Ethereum [53], a versatile blockhain that enables smart contracts —self-executing programs that automatically enforce the terms of an agreement. These provide the underlying infrastructure for non-fungible tokens (NFTs), unique identifiers often used to represent ownership of digital objects (e.g, art and collectibles). Unlike fungible cryptocurrencies, NFTs are unique and non-interchangeable. Smart contracts also enable decentralized autonomous organizations (DAOs), a form of token-based governance without a central authority, allowing for distributed decision-making and management [28].
B.   Decentraland
@@
Publicly released in February 2020, Decentraland’s social virtual world revolves around the so-called tokenomics, i.e., an economy based on the exchange of assets via cryptographic tokens on blockchains [36]. Decentraland has its own cryptocurrency, MANA, with which users can buy in-world NFTs for digital assets such as land parcels and wearables. Possession of MANA is also used as the basis of voting power for platform decisions via Decentraland’s DAO [22].
@@
The world (see Figure II) is composed of 90,601 square land parcels distributed on a 301$\times$301 grid. Before the public release of the project, 37.4% of the world map was allocated for districts, i.e., groups of adjacent parcels of privately owned land that share similar user interests. Among the biggest districts we find, for example, the gambling-focused Vegas City and the shopping-oriented Fashion Street.
@@
Figure II: Decentraland world map, where the biggest districts are indicated in violet and roads in gray. District Vegas City is home to Diamond Hands City and The Stronghold (both highlighted in yellow), ICE Poker casinos owned by Decentral Games. They cover only a minuscule surface of the map: less than 0.1%.
@@
Decentraland wearables are virtual garments, accessories or full body costumes (called skins) used to change the appearance of one’s avatar. Wearables are grouped into collections of one or more items. Each item represents a distinct design from which one or more NFTs can be minted, i.e, a new unique token ID representing the asset is created and assigned. In addition, a given item has a rarity limit (from a predefined list) that sets its maximum supply of tokens —e.g., a unique item can be minted only once, while a common item 100,000 times. To publish a wearable collection, creators pay a fee and submit their designs to the approval committee appointed by the DAO. Once approved, creators can sell their wearables via Decentraland’s Marketplace (for a fee that goes to the DAO), or they can mint and transfer the respective NFTs to a blockchain address (their own or not), either as a gift or to be sold on a different marketplace.
C.   Decentral Games
@@
The project DG started in 2019 and has played a part in many initiatives in Decentraland ever since. In fact, DG established one of the first in-world casinos, with games such as roulette or poker, and it was a prominent creator of wearables since the beginning. DG also established its own DAO, as well as its utility token and namesake DG, which grants access to rewards and voting power in the project’s governance. In October 2021, DG released ICE Poker, touted to be a Metaverse Play-to-Earn (P2E) game, i.e., one in which users are able to earn cryptocurrency simply by playing. In preparation, DG also launched the cryptocurrency ICE to function as the in-game currency for rewards and minting wearables. The following month, DG also published several exclusive ICE-themed wearables that granted attractive bonuses while playing the game, based on their ranking level.
@@
Inside Decentraland, there are two DG virtual casinos dedicated to ICE Poker, both located in the Vegas City district. The first and more popular is called The Stronghold; the second, Diamond Hands City, is exclusive to ICE-themed wearables with the highest rank. Outside Decentraland, DG wearables also give access to a related but separate mobile and web game released on September 2022, called ICE Poker Arcade. In this version of the game, instead of cryptocurrency players earn electronic gadgets, DG merchandise or other virtual assets.
@@
The level of access and/or reward provided by DG wearables varies by item, as illustrated by the examples in Figure III; the more expensive the better. Within either game mode, players can wear up to five wearables to augment their rewards. For certain ICE-themed wearables, it is also possible to upgrade their ranking level (ranging from 1 to 5) to gain better rewards by paying a smaller fee compared to buying one. Additionally, all DG wearables can be delegated to other players, who can then split their earnings with the wearable’s owner, with the latter retaining the rights over the asset on the blockchain and being able of revoking the delegation at any moment.
@@
Figure III: Examples of wearables created by Decentral Games: a) Venetian Mask (ICE Rank 1), which gives 1–7% bonus; b) DG Suited Tracksuit Bottom, which can be used to play Arcade Mode; and c) Paludamentum (ICE Rank 5), which gives 35–45% bonus.
III.   Related Work
@@
While we are still in the early development stages of the Metaverse, numerous independent virtual worlds have already sprung up over the past few decades. Notable examples include Second Life, Meta Horizon Worlds, and massively multiplayer online (MMO) games such as World of Warcraft, Minecraft, Roblox and Fortnite. However, these virtual environments have faced significant criticism regarding the governance and ownership of in-world assets. In Second Life, for instance, there has been a perceived emergence of virtual feudalism [26]. Roblox has been criticized for profiting from child labor [45] and for allowing harmful content and games with gambling-like mechanisms in its user-generated virtual worlds [35]. World of Warcraft has also grappled with conflicts between the game’s publisher and players over agency and ownership of in-game assets [20]. More generally, there has been a separation between content and platform ownership in virtual worlds, often to the detriment of users, whose rights over bought or created in-platform content are usually restricted [58].
@@
In response, a wave of blockchain-based platforms has emerged [9]. Virtual worlds such as Decentraland, The Sandbox, Axie Infinity, and Upland, among others, attempt to mitigate the aforementioned issues by leveraging blockchain’s transparency, security, and decentralization for digital ownership, cryptocurrency transactions, and governance. Decentraland [43] is a standout example, being the earliest and most popular virtual world based on blockchain. It has reached a wide audience and gone through many changes over several years, hence it offers a trove of publicly available data for the study of emerging phenomena on these novel asset-oriented platforms. However, existing research on Decentraland has primarily focused on landholding assets [14, 21, 27] or examined the platform’s tokens within the broader context of the NFT ecosystem [39]. There is a lack of studies regarding other kinds of digital assets available in Decentraland, namely avatar wearables, with the latter having attracted the attention of the fashion industry [24].
A.   Fashion in Virtual Worlds